Suppose you have 50 kg of a substance that has a half-life of five years. How much would remain after twenty years?
Answer by ankor@dixie-net.com(22740)

Suppose you have 50 kg of a substance that has a half-life of five years.
How much would remain after twenty years?
:
The radioactive decay formula: A = Ao*2^(-t/h), where
A = amt after t time
Ao = initial amt
t = time of decay
h = half-life of substance
:
A = 50*2^(-20/5)
A = 50*2^-4
A = 50*.0625
A = 3.125 kg after 20 yrs